# SCKansformer: Fine-Grained Classification of Bone Marrow Cells via Kansformer Backbone and Hierarchical Attention Mechanisms

### v1

@@ 3.9.1 Comparison of Model Performance
As shown in Table 1, we further validate the sophistication of the SCKansformer model by experimentally comparing it with other typical classification models and advanced microscopic image cell classification models on the BMCD-FGCD dataset. Compared with other typical or state-of-the-art models, SCKansformer exhibits significant advantages in handling the microscopic image classification task.
Compared to the VGG family of models, although these classical models perform consistently in traditional image classification tasks, their network depth results in significant computational overhead and memory consumption. These issues become particularly pronounced when dealing with high-resolution microscopic images. In contrast, the SCKansformer adopts a more efficient KAN network in place of the traditional MLP layer, thereby reducing memory consumption. Additionally, through its global-local attention mechanism, the SCKansformer excels at capturing subtle features in microscopic images. This results in improvements in F1-score and Accuracy, outperforming the highest values of the VGG models by 9.79 and 8.73 percentage points, respectively.
In comparison with the ResNet series models, the SCKansformer also demonstrates exceptional performance in processing microscopic images. The ResNet series models achieve a highest F1-score of 75.55, whereas the SCKansformer reaches an impressive 84.34, marking an improvement of 8.79 percentage points. In terms of Accuracy, the best performance of the ResNet series is 75.48, while the SCKansformer model achieves 83.23, an enhancement of 7.75 percentage points. By combining MSA and Local Part module, the SCKansformer effectively captures both global and local features of microscopic images, thereby significantly improving classification accuracy.
@@
Method Precision Recall F1-score Accuracy
VGG16[23] 74.68 74.43 74.55 74.50
VGG19[23] 73.53 73.35 73.44 73.26
ResNet-50[24] 75.62 75.49 75.55 75.48
ResNet-101[24] 75.43 75.20 75.31 75.21
WBCsNet[27] 73.65 72.04 72.84 72.43
BloodCaps[28] 78.72 76.74 77.94 78.31
ViT[25] 77.65 77.41 77.53 77.42
EfficientV2[26] 77.37 77.21 77.29 77.20
WBC-GLAformer[30] 79.82 79.38 79.60 79.45
WBC YOLO-ViT[29] 78.92 79.31 79.67 78.47
SCKansformer(Ours) 85.82 84.28 84.34 83.23
@@ 3.9.2 Validation of Model Generalizability
To further validate the generalizability of our proposed SCKansformer model, we compare it with other typical classification models and advanced microimaging cell classification models on two publicly available leukocyte datasets, PBC and ALL-IDB. As shown in Table 2 and Table 3, the SCKansformer model exhibits strong generalizability and achieves superior results on both datasets.
To better visualize the misclassified cells, we also plot the confusion matrix in this paper. As illustrated in Fig. 6(a), within the PBC dataset, similarly, in Fig. 6(b) illustrated that in the ALL-IDB dataset, only very few leukemia cells are misclassified, and the misclassification rate is very low. The generalizability validation further demonstrates that our proposed SCKansformer model performs well on cellular imaging datasets with uneven categories and adapts effectively to datasets with balanced categories, such as ALL-IDB and PBC, yielding promising results.
@@
Method Precision Recall F1-score Accuracy
VGG16[23] 98.02 98.01 98.01 98.01
VGG19[23] 97.31 97.28 97.29 97.28
ResNet-50[24] 98.83 98.83 98.83 98.83
ResNet-101[24] 98.83 98.83 98.83 98.83
WBCsNet[27] 97.34 97.33 97.33 97.33
BloodCaps[28] 98.80 97.77 97.77 97.77
ViT[25] 94.60 94.62 94.58 94.62
EfficientV2[26] 98.90 98.89 98.89 98.89
WBC-GLAformer[30] 99.40 99.39 99.39 99.40
WBC YOLO-ViT[29] 98.83 98.82 98.82 98.82
SCKansformer(Ours) 99.47 99.46 99.46 99.46
@@
Method Precision Recall F1-score Accuracy
VGG16[23] 96.43 96.15 96.15 96.15
VGG19[23] 98.15 98.08 98.08 98.08
ResNet-50[24] 96.15 96.15 96.15 96.15
ResNet-101[24] 96.15 96.15 96.15 96.15
WBCsNet[27] 95.87 95.69 95.69 95.69
BloodCaps[28] 97.32 97.31 97.31 97.31
ViT[25] 98.15 98.08 98.08 98.08
EfficientV2[26] 98.15 98.08 98.08 98.08
WBC-GLAformer[30] 99.15 99.10 99.09 99.10
WBC YOLO-ViT[29] 97.32 97.27 97.27 97.28
SCKansformer(Ours) 99.90 99.85 99.85 99.86
Figure 6: Confusion matrix for public datasets PBC and ALL-IDB. (a) is the confusion matrix of PBC dataset. (b) is the confusion matrix of ALL-IDB dataset.
@@ 3.10.1 Ablation of Model Performance
In the ablation experiment section, we further analyze the effects of the Kansformer Encoder, SCConv Encoder, and GLAE components on the classification performance of the SCKansformer model across three datasets. We evaluate the performance of the SCKansformer model by comparing the removal of the Kansformer Encoder, SCConv Encoder, and GLAE components, as well as the performance of the complete SCKansformer model. This evaluation is conducted using Precision, Recall, F1-score, and Accuracy metrics to determine the impact of each module on the overall model. As illustrated in Table 4, the results indicate that all three modules positively influence the model’s classification performance. Specifically:
@@
Method Precision Recall F1-score Accuracy
SCKansformer (w/o SCConv Encoder) 83.79 82.71 82.14 81.77
SCKansformer (w/o Global-Local Attention Encoder) 75.38 75.31 74.79 74.87
SCKansformer (w/o Kansformer Encoder KAN layer) 83.27 81.23 81.34 82.37
SCKansformer (Ours) 85.82 84.28 84.34 83.23
@@ 3.10.2 Validation of Model Generalizability
Additionally, we conducted the same ablation experiments on the publicly available leukocyte datasets, PBC and ALL-IDB, which feature balanced categories. These experiments further confirmed the significant impact of each key component on classification accuracy. As shown in Table 5 and Table 6, the removal of any component led to a marked decline in classification accuracy, emphasizing the importance of these components in enhancing the model’s feature representation capabilities, optimizing feature expression, and effectively capturing both global and local features of microscopic images.
Combined with the ablation experiment results on the imbalanced BMCD-FGCD dataset, these supplementary experiments provide comprehensive support for the robustness and generalizability of the SCKansformer model.
@@
Method Precision Recall F1-score Accuracy
SCKansformer (w/o SCConv Encoder) 97.34 97.31 97.31 97.32
SCKansformer (w/o Global-Local Attention Encoder) 88.01 88.01 87.93 88.03
SCKansformer (w/o Kansformer Encoder KAN layer) 95.13 95.14 95.14 95.13
SCKansformer (Ours) 99.47 99.46 99.46 99.46
@@
Method Precision Recall F1-score Accuracy
SCKansformer (w/o SCConv Encoder) 96.43 96.21 96.12 96.23
SCKansformer (w/o Global-Local Attention Encoder) 89.43 86.51 86.31 86.53
SCKansformer (w/o Kansformer Encoder KAN layer) 98.17 98.12 98.14 98.13
SCKansformer (Ours) 99.90 99.85 99.85 99.86
Conclusion
This paper proposes a novel fine-grained classification model, SCKansformer, for bone marrow blood cells, addressing the issues of insufficient feature expressiveness, poor interpretability, and feature redundancy in current bone marrow blood cell classification methods. The network structure primarily includes the Kansformer Encoder, SCConv Encoder, and GLAE components. The Kansformer Encoder utilizes the KAN network in place of the traditional MLP layer. Through the KAN network, the model assigns learnable activation functions to the edge weights, significantly enhancing the model’s nonlinear feature expression capability and interpretability. This enables the model to accurately extract and represent complex cellular features. The SCConv Encoder optimizes feature representation through the SRU and CRU. It effectively eliminates complex and confusing background information in microscopic images, reduces feature redundancy, and ensures the model focuses on key cellular features. The GLAE component captures both global and local features of microscopic images by integrating a MSA module with a Local Part module. This combination addresses the difficulty of simultaneously capturing global and local features, which often involve long-distance dependencies in cellular features. Consequently, it greatly improves the accuracy and reliability of classification. Through these innovative designs, the SCKansformer model excels in managing the complexity and accuracy requirements of myeloid blood cell classification, providing an efficient and reliable solution for microscopic image analysis.
In addition, the development of myeloid blood cell classification technology has been significantly constrained by the size and quality of existing datasets. To address this, we have jointly developed the BMCD-FGCD dataset with the Department of Hematology at Zhejiang Hospital. This dataset contains over 10,000 samples with nearly 40 classifications of highly fine-grained annotations. We have chosen to make this private dataset publicly available to the academic community. The establishment and disclosure of the BMCD-FGCD dataset not only provide a robust database for our research but also serve as an important resource for advancing the field. We anticipate that the release of this high-quality private dataset will promote further development in the field of myeloid hematopoietic cell classification.
@@ Conclusion
In comparative experiments, we evaluate the SCKansformer model against other advanced cell classification models using three datasets: the private BMCD-FGCD dataset and the publicly available PBC and ALL-IDB datasets. The experimental results demonstrate that SCKansformer significantly outperforms existing classification methods and advanced microcell classification methods across these datasets, excelling in metrics such as Accuracy, Precision, Recall, and F1-score. Additionally, we conduct ablation experiments on the BMCD-FGCD, PBC, and ALL-IDB datasets to determine the importance of key components of the model (e.g., the Kansformer Encoder, SCConv Encoder, and GLAE). The results validate the positive impact of these components on the overall model performance, effectively addressing the issues of low feature expressiveness, poor interpretability, and redundancy in cellular feature extraction encountered in existing automated methods for high-dimensional microscopic image data.
While our study significantly contributes to the field, we acknowledge its limitations. To enhance the robustness and generalizability of the SCKansformer model, future research should focus on accumulating larger and more diverse datasets for further validation. Moreover, given the rapid advancements in medical imaging technologies and deep learning methods, continuous refinement and adaptation of our model are essential to maintain its relevance and utility in real-world applications. Subsequently, we will continue to improve our BMCD-FGCD dataset and continue to publicize subsequent improved versions to further advance the field.

### v2

@@ 4.5.1 Comparison of Model Performance
As shown in Table 1, we further validate our model by experimentally comparing it with other typical classification models and advanced microscopic image cell classification models on the BMCD-FGCD dataset. Compared with other typical or state-of-the-art models, SCKansformer exhibits significant advantages in handling the microscopic image classification task.
@@
Method Precision Recall F1-score Accuracy Balanced Accuracy
VGG16 [21] 74.68 74.43 74.55 74.50 35.79
VGG19 [21] 73.53 73.35 73.44 73.26 37.63
ResNet-50 [22] 75.62 75.49 75.55 75.48 49.76
ResNet-101 [22] 75.43 75.20 75.31 75.21 50.50
WBCsNet [25] 73.65 72.04 72.84 72.43 47.50
BloodCaps [26] 78.72 76.74 77.94 78.31 54.31
ViT [23] 77.65 77.41 77.53 77.42 70.47
EfficientV2 [24] 77.37 77.21 77.29 77.20 71.51
WBC-GLAformer [28] 79.82 79.38 79.60 79.45 74.31
WBC YOLO-ViT [27] 78.92 79.31 79.67 78.47 77.41
SCKansformer (Ours) 85.82 84.28 84.34 83.23 82.91
@@ 4.5.2 Validation of Model Generalizability
To further validate the generalizability of our proposed SCKansformer model, we compare it with other typical classification models and advanced microimaging cell classification models on two publicly available leukocyte datasets, PBC and ALL-IDB. As shown in Table 2 and Table 3, the SCKansformer model exhibits strong generalizability and achieves superior results on both datasets. The generalizability validation further demonstrates that our proposed SCKansformer model performs well on cellular imaging datasets with uneven categories and adapts effectively to datasets with balanced categories, such as ALL-IDB and PBC, yielding promising results.
@@
Method Precision Recall F1-score Accuracy Balanced Accuracy
VGG16 [21] 98.02 98.01 98.01 98.01 98.30
VGG19 [21] 97.31 97.28 97.29 97.28 97.25
ResNet-50 [22] 98.83 98.83 98.83 98.83 98.99
ResNet-101 [22] 98.83 98.83 98.83 98.83 98.99
WBCsNet [25] 97.34 97.33 97.33 97.33 96.93
BloodCaps [26] 98.80 97.77 97.77 97.77 97.25
ViT [23] 96.90 96.90 96.89 96.90 96.75
EfficientV2 [24] 98.74 98.74 98.74 98.74 98.84
WBC-GLAformer [28] 99.40 99.39 99.39 99.40 98.79
WBC YOLO-ViT [27] 98.83 98.82 98.82 98.82 98.31
SCKansformer (Ours) 99.47 99.46 99.46 99.46 99.37
@@
Method Precision Recall F1-score Accuracy MCC
VGG16 [21] 96.43 96.15 96.15 96.15 92.58
VGG19 [21] 98.15 98.08 98.08 98.08 89.06
ResNet-50 [22] 96.15 96.15 96.15 96.15 92.58
ResNet-101 [22] 96.15 96.15 96.15 96.15 88.53
WBCsNet [25] 95.87 95.69 95.69 95.69 93.25
BloodCaps [26] 97.32 97.31 97.31 97.31 96.44
ViT [23] 98.15 98.08 98.08 98.08 96.22
EfficientV2 [24] 98.15 98.08 98.08 98.08 96.23
WBC-GLAformer [28] 99.15 99.10 99.09 99.10 97.52
WBC YOLO-ViT [27] 97.32 97.27 97.27 97.28 96.21
SCKansformer (Ours) 99.90 99.85 99.85 99.86 99.83
@@ 4.6.1 Ablation of Model Performance
In the ablation experiment section, we further analyze the effects of the Kansformer Encoder, SCConv Encoder, and GLAE components on the classification performance of the SCKansformer model across three datasets. We evaluate the performance of the SCKansformer model by comparing the removal of the Kansformer Encoder, SCConv Encoder, and GLAE components, as well as the performance of the complete SCKansformer model. As illustrated in Table 4, the results indicate that all three modules positively influence the model’s classification performance. Specifically:
@@
Method Precision Recall F1-score Accuracy Balanced Accuracy
w/o SCConv Encoder 83.79 82.71 81.14 81.77 77.50
w/o GLAE 75.38 75.31 74.79 74.87 70.33
w/o Kansformer Encoder KAN layer 82.27 81.23 81.34 82.37 80.47
SCKansformer (Ours) 85.82 84.28 84.34 83.23 82.91
@@ 4.6.2 Validation of Model Generalizability
Additionally, we conducted the same ablation experiments on the publicly available leukocyte datasets, PBC and ALL-IDB, which feature balanced categories. These experiments further confirmed the significant impact of each key component on classification accuracy. As shown in Table 5 and Table 6, the removal of any component led to a marked decline in classification accuracy, emphasizing the importance of these components in enhancing the model’s feature representation capabilities, optimizing feature expression, and effectively capturing both global and local features of microscopic images. Combined with the ablation experiment results on the imbalanced BMCD-FGCD dataset, together they provide comprehensive support for the robustness and generalizability of the SCKansformer model.
@@
Method Precision Recall F1-score Accuracy Balanced Accuracy
w/o SCConv Encoder 97.34 97.31 97.31 97.32 97.36
w/o GLAE 88.01 87.99 87.87 87.99 84.97
w/o Kansformer Encoder KAN layer 95.13 95.14 95.14 95.13 92.70
SCKansformer (Ours) 99.47 99.46 99.46 99.46 99.37
@@
Method Precision Recall F1-score Accuracy MCC
w/o SCConv Encoder 96.43 96.15 96.12 96.15 92.58
w/o GLAE 89.43 86.51 86.29 86.53 75.88
w/o Kansformer Encoder KAN layer 98.15 98.08 98.08 98.08 96.23
SCKansformer (Ours) 99.90 99.85 99.85 99.86 99.83
5 Conclusion
@@
In this paper, we propose SCKansformer, a novel fine-grained classification model for bone marrow blood cells that addresses key challenges such as insufficient feature expressiveness, poor interpretability, and feature redundancy in current methods. We evaluate the SCKansformer model against other advanced cell classification models using three datasets: the private BMCD-FGCD dataset and the publicly available PBC and ALL-IDB datasets. The experimental results show that SCKansformer significantly outperforms existing advanced microcell classification methods across these datasets, excelling in all metrics. Additionally, we conduct ablation experiments on the BMCD-FGCD, PBC, and ALL-IDB datasets to determine the importance of key components of the model. The results validate the positive impact of these components on the overall model performance.
In addition, the development of myeloid blood cell classification technology has been significantly constrained by the size and quality of existing datasets. To address this issue, we have collaborated with the Department of Hematology at Zhejiang Hospital to develop the BMCD-FGCD dataset. This dataset contains over 10,000 samples with nearly 40 fine-grained classifications. We have decided to make this private dataset publicly available to the academic community. The establishment and release of the BMCD-FGCD dataset not only provide a robust database for our research but also serve as a valuable resource for advancing the field. We anticipate that the availability of this high-quality dataset will promote further development in myeloid blood cell classification. Future work will focus on expanding dataset diversity and refining the model to ensure its applicability in clinical settings.